# IPAD: Inverse Prompt for AI Detection - A Robust and Explainable LLM-Generated Text Detector

### v1

@@ 3.1.2 Robustness across different LLMs
The results of IPAD for detecting the dataset OUTFOX Koike et al. (2024) across LLMs are presented in Table 1 and Table 2, respectively. They show that both versions are highly robust across various LLMs, while  Regeneration Comparator is a bit more efficient.
As for  Regeneration Comparator, when the original generator and re-generator are the same model, the performance is optimal. However, even when the re-generator is different from the original generator, the results remain impressive with ChatGPT used as the re-generator. These results imply that, in practical applications, it is possible to use a common set of LLMs as re-generators. If one or more correponding  Distinguishers from different LLMs classify the results as ’yes’, it can be inferred that the text is likely to be LGT, whereas if all  Distinguishers classify the results as ’no’, the text is more likely to be HWT. Furthermore, for applications aiming to save computational resources and improve efficiency, using ChatGPT as the sole re-generator still yields robust performance across all tested models.
Original Generator Metrics (%)
HumanRec MachineRec AvgRec F1
ChatGPT 98.00% 99.80% 98.90% 98.89%
GPT-3.5 97.20% 99.90% 98.55% 98.53%
Qwen-turbo 98.00% 98.10% 98.05% 98.05%
Llama-3-70B 98.00% 100.00% 99.00% 98.99%
Table 1: IPAD with  Prompt-Text Consistency Verifier performance on different LLMs
Original Generator Re-Generator Metrics (%)
HumanRec MachineRec AvgRec F1
ChatGPT ChatGPT 99.70% 100.00% 99.85% 99.85%
GPT-3.5 GPT-3.5 98.00% 100.00%  99.00%  99.00%
ChatGPT 97.00% 100.00% 98.50% 98.50%
Qwen-turbo Qwen-turbo 98.00% 98.40%  98.20%  98.20%
ChatGPT 99.70% 94.40% 97.05% 97.13%
Llama-3-70B Llama-3-70B 96.60% 100.00% 98.30% 98.30%
ChatGPT 99.70% 99.40%  99.55%  99.55%
Table 2: IPAD with  Regeneration Comparator performance on different LLMs
@@ 3.1.3 Comparison of IPAD with other detectors in and out of distribution
Table 3 compares the performance of two versions of IPAD with other detection methods in the OUTFOX dataset with and without attacks Koike et al. (2024). The results show that both versions of IPAD generally outperform other detectors, while that IPAD with  Prompt-Text Consistency Verifier for detecting ChatGPT with DIPPER attack performs worse. These results imply that IPAD with  Regeneration Comparator demonstrates superior robustness compared to alternative detection methods in the OUTFOX dataset with and without attacks.
Original Generator Detection Methods Metrics (%)
HumanRec MachineRec AvgRec F1
ChatGPT RoBERTa-base 93.80% 92.20% 93.00% 92.90%
RoBERTa-large 91.60% 90.00% 90.80% 90.70%
HC3 detector 79.20% 70.60% 74.90% 73.80%
OUTFOX 97.80% 92.40% 95.10% 95.00%
IPAD version1 98.00% 99.80% 98.90% 98.89%
IPAD version2 99.70% 100.00% 99.85% 99.85%
GPT-3.5 RoBERTa-base 93.80% 92.00% 92.90% 92.80%
RoBERTa-large 92.60% 92.00% 92.30% 92.30%
HC3 detector 79.20% 85.00% 82.10% 82.60%
OUTFOX 97.60% 96.20% 96.90% 96.90%
IPAD version1 97.20% 99.90% 98.55% 98.53%
IPAD version2 97.00% 100.00% 98.50% 98.50%
ChatGPT with DIPPER Attack RoBERTa-base 93.80% 89.20% 91.50% 91.30%
RoBERTa-large 91.60% 97.00% 94.30% 94.40%
HC3 detector 79.20% 3.40% 41.30% 5.50%
OUTFOX 98.60% 66.20% 82.40% 79.00%
IPAD version1 98.00% 75.10% 86.55% 87.93%
IPAD version2 99.70% 95.40% 97.55% 97.60%
ChatGPT with OUTFOX Attack RoBERTa-base 93.80% 69.20% 81.50% 78.90%
RoBERTa-large 91.60% 56.20% 73.90% 68.30%
HC3 detector 79.20% 0.40% 39.80% 0.70%
OUTFOX 98.80% 24.80% 61.80% 39.40%
IPAD version1 98.00% 95.40% 96.70% 96.74%
IPAD version2 99.70% 98.00% 98.85% 98.86%
Table 3: Comparison of IPAD with other detectors on in-distribution data, where  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator
Table 4 presents the performance of various detection methods on OOD datasets to assess their generalizability, where the baseline data refer to DetectRL  Wu et al. (2024). The results demonstrate that IPAD with  Regeneration Comparator consistently outperforms all other baselines in all OOD datasets with and without attacks. In contrast, IPAD with  Prompt-Text Consistency Verifier exhibits strong performance on OOD datasets without attacks but shows a noticeable drop in effectiveness when subjected to attacks. For instance, while it achieves competitive results on datasets like XSum (99.90%) and Writing (99.20%), its performance against attacks, such as Prompt Attack (86.90%) and Paraphrase Attack (82.72%), is significantly lower than IPAD with  Regeneration Comparator. This suggests that IPAD with  Regeneration Comparator demonstrates better generalizability and robustness.
OOD Datasets or attack type Detection Methods
LRR Fast-DetectGPT Rob-Base IPAD with version1 IPAD version2
Arxiv 48.17% 42.00% 81.06% 84.47% 98.60%
XSum 48.41% 45.72% 76.81% 99.90% 98.90%
Writing 58.70% 51.13% 86.29% 99.20% 95.80%
Review 58.21% 54.55% 87.84% 98.50% 89.30%
Avg. for non-attacked datasets 53.37% 48.35% 83.00% 95.52% 95.65%
Prompt Attack 54.97% 43.89% 92.81% 86.90% 93.05%
Paraphrase Attack 49.23% 41.15% 90.02% 82.72% 95.89%
Perturbation Attack 53.62% 44.38% 92.12% 94.96% 95.32%
Avg. for attacked datasets 52.61% 43.14% 91.65% 88.26% 94.75%
Avg. 53.04% 46.12% 86.70% 92.41% 95.26%
Table 4: The performance of IPAD in generalization assessment (AUROC). The selected detectors are evaluated on OOD data, all sourced from and processed using the DetectRL baseline, where  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator.
3.1.4 Robustness conclusion
Our experimental results demonstrate that both IPAD versions exhibit strong performance across different LLMs, outperforming existing detection methods and maintaining robustness on OOD datasets. The IPAD with  Regeneration Comparator outperforming baselines by 9.73% (F1-score) on in-distribution data and 12.65% (AUROC) OOD data. Notably, IPAD with  Regeneration Comparator achieves significantly better performance than IPAD with  Prompt-Text Consistency Verifier in attack scenarios of 3.78% (F1-score). While IPAD with  Prompt-Text Consistency Verifier performs robustly in standard settings, its performance declines when facing attacks. The calculation of these statistics are shown in Appendix  B.
3.2 Necessity and Effectiveness of Prompt Inverter and Distinguishers
3.2.1 Necissity of the Prompt Inverter and Distinguishers
@@
To prove that it is necessary to fine-tune on IPAD with IPAD with  Prompt-Text Consistency Verifier and  Regeneration Comparator, we conducted ablation study to use the same finetune method on only  input texts and only  predicted prompts. The instructions are  "Is this text generated by LLM?", and  "Prompt Inverter predicts prompt that could have generated the input texts. Is this prompt predicted by an input texts written by LLM?", respectively.
The results shown in Figure  2 from the ablation study show that fine-tuning on either only the  input text or only the  predicted prompt leads to poor performance. This underscores the importance of fine-tuning on a combination of both the input text and predicted prompt, as explored in the  Prompt-Text Consistency Verifier, or on the input text and regenerated text, as examined in the  Regeneration Comparator, for more effective detection.
Figure 2: Ablation Study Results. The  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator.
3.2.2 The effectivenss of the IPAD Prompt Inverter
We use DPIC Yu et al. (2024) and PE Zhang et al. (2024c) as baseline methods for prompt extraction. DPIC employs a zero-shot approach using the prompt states in Appendix  C, while PE uses adversarial attacks to recover system prompts.
@@ 3.2.2 The effectivenss of the IPAD Prompt Inverter
In our evaluation, we tested 1000 LGT and 1000 HWT samples. We use only in-distribution data for testing since only these datasets include original prompts. The metrics are all tested on comparing the similarity of the original prompts and the predicted prompts. The results shown in Table  5 illustrate that IPAD consistently outperforms both DPIC and PE across all four metrics (BartScore Yuan et al. (2021), Sentence-Bert Cosine Similarity Reimers and Gurevych (2019), BLEU Papineni et al. (2002), and ROUGE-1 Lin (2004)), which highlight the effectiveness of the IPAD  Prompt Inverter.
Evaluation Bart-large-cnn Sentence-Bert BLEU ROUGE-1
LGT
DPIC -2.12 0.46 5.61E-05 0.04
PE -2.23 0.58 3.21E-04 0.25
IPAD  -1.84  0.69  0.24  0.51
HWT
DPIC -2.47 0.42 8.75E-06 0.06
PE -2.39 0.53 2.56E-08 0.13
IPAD  -2.22  0.57  1.30E-01  0.39
Table 5: Comparison of the IPAD Prompt Inverter with other prompt extractors
3.2.3 The Effectiveness of the IPAD Distinguishers
@@
To examine the effectiveness of the IPAD  Distinguishers, we conducted a comparison study using the same dataset but different distinguishing methods. The first and second methods employed Sentence-Bert  Reimers and Gurevych (2019) and Bart-large-cnn  Yuan et al. (2021) to compute the similarity score between the input texts and the regenerated texts. We selected thresholds that maximized AvgRec, which were 0.67 for Sentence-Bert and -2.52 for Bart-large-cnn. The classification rule is that the texts with scores greater than the threshold will be classified as LGT, while the texts with scores less than or equal to the threshold will be classified as HWT.
The third and fourth methods involved directly prompting ChatGPT as follows:
Instruction:  "Text 1 is generated by an LLM. Determine whether Text 2 is also generated by an LLM with a similar prompt. Answer with only YES or NO."  Input:  "Text 1: {Regenerated Text}; Text 2: {LGT} or {HWT}".
and  Instruction:  "Can LLM generate text2 through the prompt text1? Answer with only YES or NO." with  Input:  "Text 1: {Predicted Prompt}; Text 2: {Input text}".
The final results demonstrated that the other distinguishing methods performed worse than the two IPAD  Distinguishers, highlighting the superior effectiveness of the IPAD  Distinguishers.
Distinguish Method HumanRec MachineRec AvgRec F1
Sentence-Bert (Threshold 0.67) 61.20% 95.20% 78.20% 63.51%
Bart-large-cnn (Threshold -2.52) 42.60% 97.20% 69.90% 43.96%
Prompt to ChatGPT version 1 33.20% 64.50% 48.85% 44.77%
Prompt to ChatGPT version 2 12.50% 100% 56.25% 12.50%
IPAD version 1 98.00% 99.80% 98.90% 98.10%
IPAD version 2  99.70%  100%  99.85%  99.70%
Table 6: Comparison of Different Distinguishers, where  IPAD version1 stands for  IPAD with  Prompt-Text Consistency Verifier and  IPAD version2 stands for  IPAD with  Regeneration Comparator.
3.3 Explanability Assessment of IPAD
3.3.1 Different Linguistic Features of HWT prompts and LGT prompts
One of the primary distinctions between the HWT prompts and the LGT prompts is sentence complexity. LGT prompts are typically more complex, characterized by longer sentence lengths (mean value of 1.514 and 1.794), higher syllable counts (mean values of total syllabus three are 1.572 and 3.042), and more stop-words (mean values of 9.88 and 10.045). HWT prompts, on the other hand, are characterized by shorter, less complex sentences that are easier to process and understand, as examples shown in Appendix D Figure 3.
Beyond the differences in syntactic complexity, we also explored variables in LIWC. We did the difference comparison by using HWT and LGT prompts as inputs for ChatGPT, for example, instructing with the prompts ’determine the pronoun usage of this sentence, answer first person, second person, or third person’ and ’determine the level of analytical thinking of these sentences, answer a number from 1 to 5’. The results show that there are distinguish difference in pronoun usage and analytical thinking level. The HWT prompts frequently use second-person pronouns (e.g., ’you’) - 75 occurrences per 1,000 prompts - due to the subjective tone often employed in HWT. In contrast, LGT prompts primarily feature first- and third-person pronouns, with second-person pronouns appearing only 2 per 1,000 prompts. LGT prompts typically present instructions and questions in a more objective manner. As shown in Appendix  D Figure  4, LGT prompts show higher analytical thinking levels than HWT prompts. With level 1 as the lowest and level 5 as the highest, LGT has 68.9% of level 4 and 24.3% of level 5, but HWT has only 48.0% of level 4, and 0.8% of level 5. It suggests that LGT prompts encourage more analytical thinking, while HWT prompts tend to focus more on concrete examples, with less emphasis on critical analysis, as examples shown in Appendix D Figure  5.
3.4 User Study
@@
To assess the explainability improvement of IPAD, we designed an IRB-approved user study with ten participants evaluating one HWT and one LGT article. We used IPAD version 2 due to its superior OOD performance and attack resistance. Participants compared three online detection platforms with screenshots shown in Appendix E [^5] https://www.scribbr.com/ai-detector/ [^6] https://quillbot.com/ai-content-detector [^7] https://app.gptzero.me/ with IPAD’s process (which displayed input texts, predicted prompts, regenerated texts, and final judgments). After evaluation, users rated IPAD on four key explainability dimensions. Transparency received strong ratings (40%:5, 60%:4), with users appreciating the visibility of intermediate processes. Trust scores were more varied (10%:3, 70%:4, 20%:5), but IPAD was generally considered more convincing than single-score detectors. Satisfaction was mixed (30%:3, 30%:4, 40%:5), with users acknowledging better detection but raising concerns about energy efficiency since IPAD runs three LLMs. Debugging received unanimous 5s, as users could easily analyze the predicted prompt and regenerated text to verify the decision-making process. If needed, users could refine the generated content by adjusting instructions, such as specifying a word count, making IPAD a more effective and user-friendly tool compared to black-box detectors.
@@ 4.1 AI detectors Methods and challenges
Watermarking technology inserts specific patterns into training datasets Shevlane et al. (2023); Gu et al. (2022) or manipulates the model output during inference to embed a watermark Lucas and Havens (2023). However, watermarking needs to access of the LLM deployment and can face attacks, such as identifying and erasing the watermark Hou et al. (2024).  Statistics-based methods analyze inherent textual features to identify language patterns Kalinichenko et al. (2003); Hamed and Wu (2023), but their effectiveness depends on corpus size and model diversity Wu et al. (2025). Some other statistical methods use n-gram probability divergence Yang et al. (2024b) or similarity between original and revised texts Mao et al. (2024); Zhu et al. (2023) while still face robustness challenges under adversarial attacks Wu et al. (2025)Neural-based methods such as RoBERTa Liu et al. (2020), Bert Devlin et al. (2019), and XLNet Yang et al. (2019) have been robust in domain-specific tasks. Adversarial learning techniques are increasingly being used Yang et al. (2024a) to increase effectiveness in attacked datasets.

### v2

@@ 3.1.2 Robustness across different LLMs
As shown in Table 1, IPAD achieves consistently strong performance across all combinations of original generators and re-generators, which shows its robustness to diverse LLM as generators. The best results are generally observed when the original generator and the re-generator are the same, while the gpt-3.5-turbo serves as an effective universal re-generator: it performs well even when the original generator differs. In real-world applications where the identity of the original generator is unknown, using gpt-3.5-turbo as a fixed re-generator provides a practical and reliable solution.
Original Generator Re-Generator HumanRec MachineRec AvgRec AUROC
gpt-3.5-turbo gpt-3.5-turbo 98.50 100 99.25 100
gpt-4 gpt-4 98.70 100 99.35 100
gpt-3.5-turbo 96.10 100 98.05 99.96
Qwen-turbo Qwen-turbo 98.60 99.80 99.20 99.96
gpt-3.5-turbo 98.40 99.50 98.95 99.86
LLaMA-3-70B LLaMA-3-70B 98.70 100 99.35 100
gpt-3.5-turbo 98.60 100 99.30 100
Table 1: Detection Accuracy (HumanRec, MachineRec, AvgRec, and AUROC %) of IPAD across Various LLMs on In-Distribution Data
@@ 3.1.3 Comparison of IPAD with other detectors in and out of distribution
In Distribution.
For the in-distribution data, as shown in Figure 2, the baseline detectors like RoBERTa, HC3, and OUTFOX perform well on standard data but degrade significantly under DIPPER and OUTFOX attacks. In contrast, IPAD maintains high accuracy across all scenarios, which surpasses the strongest baseline 9.05% in AvgRec.
Figure 2: The In-distribution data performance of IPAD and the baseline detectors. Since  Koike et al. (2024) only presents the AvgRec data for the baselines, we also calculate AvgRec data for IPAD to compare.
Out of Distribution.
Table 2 reports detection accuracy across four benchmark datasets, which shows that IPAD significantly outperforms prior baselines. Table 3 further evaluates robustness under three attack types, where IPAD again demonstrates superior resilience. Compared to the strongest baseline, IPAD achieves a 12.93% relative improvement on standard datasets in AUROC and a 5.48% improvement on attack datasets.
Method Arxiv XSum Writing Review Average
DetectLLM (LRR) 48.17 48.41 58.70 58.21 53.37
DetectLLM (NPR) 53.85 34.59 54.96 50.09 48.37
Binoculars 84.03 77.39 94.38 90.00 86.45
Fast-DetectGPT 42.00 45.72 51.13 54.55 48.35
Rob-Base 81.06 76.81 86.29 87.84 83.00
IPAD Merge 100 99.85 99.40 98.25 99.38
Table 2: Detection Accuracy (AUROC %) on four diverse OOD datasets
Method Prompt Attack Paraphrase Attack Perturbation Attack Average
DetectLLM (LRR) 54.97 49.23 53.62 52.61
DetectLLM (NPR) 77.15 56.94 6.78 46.96
Binoculars 93.45 88.34 76.89 86.23
Fast-DetectGPT 43.89 41.15 44.38 43.14
Rob-Base 92.81 90.02 92.12 91.65
IPAD 97.30 96.00 98.10 97.13
Table 3: Detection Accuracy (AUROC %) on three attacked OOD datasets
Structured Prompts.
The results are shown in Table  4, while these datasets lack HWT references and are thus only evaluated using MachineRec, the strong scores suggest that IPAD maintains robustness even on structured diverse inputs, with an improvement of 9.87% against the strongest baseline in MachineRec.
Method LongWriter Code-Feedback Math Average
DetectLLM (LRR) 32.1 29.0 30.2 30.43
DetectLLM (NPR) 41.2 45.9 56.0 47.7
Binoculars 82.1 84.6 89.4 85.4
Fast-DetectGPT 12.0 11.1 15.1 12.7
Rob-Base 81.5 89.2 82.1 84.3
IPAD 97.5 92.7 95.6 95.27
Table 4: Detection Accuracy (MachineRec %) on three structured OOD datasets
3.2 Necessity and Effectiveness of the Prompt Inverter, PTCV, and RC
3.2.1 Necessity
@@
To prove that it is necessary to fine-tune on IPAD with IPAD with  PTCV and  RC, we conducted ablation study to use the same finetune method on only  input texts and only  predicted prompts, with the finetune data format shown in Appendix  D. We only experimented on Prompt Inverter + PTCV and Prompt Inverter + RC to compare with the three-moduled IPAD.
Based on the ablation study results as shown in Figure  3, fine-tuning only on input texts or only on predicted prompts performs poorly across all datasets in AUROC scores. While using Prompt Inverter + PTCV or Prompt Inverter + RC individually significantly improves performance, neither approach consistently excels across both HWT-style and LGT-style generations. In contrast, the full IPAD framework achieves consistently high performance across all settings, which demonstrates the necessity of the Prompt Inverter, PTCV, and RC modules.
Figure 3: Ablation study. Evaluating Fine-tune only on Input, Fine-tune only on Prompt, Prompt Inverter + PTCV, Prompt Inverter + RC, and IPAD on In-distribution datasets, standard OOD datasets, and attacked OOD datasets.
Metric LGT HWT
DPIC PE IPAD DPIC PE IPAD
Bart-large-cnn -2.12 -2.23 -1.84 -2.47 -2.39 -2.22
Sentence-Bert 0.46 0.58 0.69 0.42 0.53 0.57
BLEU 5.61E-05 3.21E-04 0.24 8.75E-06 2.56E-08 0.13
ROUGE-1 0.04 0.25 0.51 0.06 0.13 0.39
Table 5: Comparison of prompt inverters on the similarities of the original prompts and the predicted prompts on LGT and HWT.
Distinguish Method HumanRec MachineRec AvgRec
Sentence-Bert (Thr. 0.67) 61.20 95.20 78.20
Bart-large-cnn (Thr. -2.52) 42.60 97.20 69.90
Prompt to ChatGPT 33.20 64.50 48.85
IPAD 98.50 100.00 99.25
Table 6: Comparison of distinguishers on HumanRec, MachineRec, and AvgRec (%).
3.2.2 Effectiveness
Prompt Inverter.
@@ 3.2.2 Effectiveness
We use DPIC (Yu et al., 2024) and PE (Zhang et al., 2024b) as baseline methods for prompt extraction. DPIC employs a zero-shot approach using the prompt states in Appendix  E, while PE uses adversarial attacks to recover system prompts. In our evaluation, we tested 1000 LGT and 1000 HWT samples. We use only in-distribution data for testing since only these datasets include original prompts. The metrics are all tested on comparing the similarity of the original prompts and the predicted prompts. The results shown in Table  5 illustrate that IPAD consistently outperforms both DPIC and PE across all four metrics (BartScore (Yuan et al., 2021), Sentence-Bert Cosine Similarity (Reimers and Gurevych, 2019), BLEU (Papineni et al., 2002), and ROUGE-1 (Lin, 2004)), which highlight the effectiveness of the IPAD  Prompt Inverter.
PTCV and RC.
@@
We conducted a comparison study using the frozen Prompt Inverter but different distinguishing methods. The first and second methods employed Sentence-Bert (Reimers and Gurevych, 2019) and Bart-large-cnn (Yuan et al., 2021) to compute the similarity score between the input texts and the regenerated texts. We selected thresholds that maximized AvgRec, which were 0.67 for Sentence-Bert and -2.52 for Bart-large-cnn. The classification rule is that the texts with scores greater than the threshold will be classified as LGT, while the texts with scores less than or equal to the threshold will be classified as HWT. The third method is to directly prompt ChatGPT in Appendix  D, which mimic the fine-tuning process of PTCV and RC. The final results shown in Table  6 demonstrate that the other distinguishing methods performed worse than IPAD, highlighting the superior effectiveness of the IPAD  Distinguishers.
Compare with DPIC.
DPIC first uses a zero-shot prompt inverter to generate prompts, then applies a Siamese encoder and classifier to measure similarity between the embeddings of the original and regenerated texts. However, the classifier’s reliance on embedding similarity is ambiguous, as similar texts may stem from different prompts. IPAD addresses this by fine-tuning directly on raw texts and reformulating the task as a logical reasoning problem as shown in the instructions of PTCV and RC. Our trained Prompt Inverter outperforms DPIC’s generic zero-shot method as shown in Table 5, and IPAD also achieves better performance than DPIC overall, as results shown in Appendix  F.
3.3 Interpretability Assessment of IPAD
@@
To assess the explainability improvement of IPAD, we designed an interpretability assessment with ten participants evaluating one HWT and one LGT article. We used IPAD version 2 due to its superior OOD performance and attack resistance. Participants compared three online detection platforms [^6] https://www.scribbr.com/ai-detector/ [^7] https://quillbot.com/ai-content-detector [^8] https://app.gptzero.me/ with IPAD’s process (which displayed input texts, predicted prompts, regenerated texts, and final judgments). After evaluation, participants rated IPAD on four key explainability dimensions. Transparency received strong ratings (40%:5, 60%:4), with participants appreciating the visibility of intermediate processes. Trust scores were more varied (10%:3, 70%:4, 20%:5), but IPAD was generally considered more convincing than single-score detectors. Satisfaction was mixed (30%:3, 30%:4, 40%:5), with participants acknowledging better detection but raising concerns about energy efficiency since IPAD runs three LLMs. Debugging received unanimous 5s, as participants could easily analyze the predicted prompt and regenerated text to verify the decision-making process. If needed, users could refine the generated content by adjusting instructions, such as specifying a word count, making IPAD a more effective and user-friendly tool compared to black-box detectors. We further analyzed the different linguistic features of HWT prompts and LGT prompts as illustrated in Appendix  G.
@@ 4.1 AI detectors Methods and challenges
Watermarking technology inserts specific patterns into training datasets Shevlane et al. (2023); Gu et al. (2022) or manipulates the model output during inference to embed a watermark Lucas and Havens (2023). However, watermarking needs to access the LLM deployment and can face attacks, such as identifying and erasing the watermark Hou et al. (2024).  Statistics-based methods analyze inherent textual features to identify language patterns Kalinichenko et al. (2003); Hamed and Wu (2023), but their effectiveness depends on corpus size and model diversity Wu et al. (2025). Some other statistical methods use n-gram probability divergence Yang et al. (2024b) or similarity between original and revised texts Mao et al. (2024); Zhu et al. (2023) while still face robustness challenges under adversarial attacks Wu et al. (2025)Neural-based methods such as RoBERTa Liu et al. (2020), Bert Devlin et al. (2019), and XLNet Yang et al. (2019) have been robust in domain-specific tasks. Adversarial learning techniques are increasingly being used Yang et al. (2024a) to increase effectiveness in attacked datasets.